# LLMEasyQuant: Scalable Quantization for Parallel and Distributed LLM Inference

### v5

@@ A.2.2 Distributed Quantization Synchronization Analysis
Since AllGather is deterministic and collective, all devices receive identical copies of the global parameters.
Step 3: Consistency Guarantee After synchronization, all devices have identical quantization parameters:
$\delta_{\ell}^{(p)}=\delta_{\ell}^{\text{global}}\quad\forall p\in\{1,2,\ldots,P\}$ (39)
$z_{\ell}^{(p)}=z_{\ell}^{\text{global}}\quad\forall p\in\{1,2,\ldots,P\}$ (40)
Step 4: Quantized Weight Consistency With identical quantization parameters, all devices produce identical quantized weights:
$\hat{W}_{\ell}^{(p)}=\text{Quantize}(W_{\ell}^{(p)},\delta_{\ell}^{\text{global}},z_{\ell}^{\text{global}})=\hat{W}_{\ell}^{\text{global}}$ (41)
This ensures model consistency across all devices, completing the proof. ∎
A.3 Advanced Visualization Analysis
This section provides comprehensive visual analysis of quantization effectiveness across different models and methods using advanced visualization techniques.
A.3.1 Radar Chart Analysis
Figure 4 presents radar charts comparing LLMEasyQuant against state-of-the-art methods across multiple performance dimensions for all modern models. The radar charts visualize five key metrics:
  • Accuracy: Inverse perplexity score (higher is better)
  • Throughput: Normalized tokens per second (higher is better)
  • Memory Efficiency: Inverse memory usage (higher is better)
  • Setup Speed: Relative setup time efficiency (higher is better)
  • Calibration Efficiency: Inverse calibration data requirements (higher is better)
The radar charts reveal that SmoothQuant consistently achieves the best overall performance across all models, with superior accuracy and throughput while maintaining competitive memory efficiency. SimQuant shows strong performance in memory efficiency and calibration requirements, making it suitable for resource-constrained environments.
Figure 4: Radar Chart Analysis: Comprehensive Performance Comparison Across Modern Models
A.3.2 3D Heatmap Visualization
Figure 5 shows a 3D heatmap visualizing the relationship between model size, quantization method, and throughput performance. The visualization reveals several key insights:
  • Model Size Scaling: Larger models show more pronounced performance differences between quantization methods
  • Method Effectiveness: SmoothQuant maintains consistent performance across all model sizes
  • Throughput Patterns: The heatmap reveals optimal quantization strategies for different model scales
The 3D visualization demonstrates that LLMEasyQuant’s SmoothQuant implementation provides the most consistent performance across the entire model size spectrum, from small models (117M parameters) to large models (14B parameters).
Figure 5: 3D Heatmap: Model Size vs Quantization Method vs Throughput Performance
A.3.3 Spindle Plot Comparison
Figure 6 displays comprehensive spindle plots comparing quantization methods across different models and metrics. The four subplots analyze:
  1. 1.
    Perplexity Comparison: Shows accuracy degradation across different models
  2. 2.
    Throughput Comparison: Demonstrates inference speed across model sizes
  3. 3.
    Memory Usage Comparison: Illustrates memory efficiency gains
  4. 4.
    Efficiency Score: Combines accuracy and throughput for overall efficiency
The spindle plots reveal that:
  • SmoothQuant achieves the best perplexity scores across all models
  • LLMEasyQuant implementations maintain competitive throughput
  • Memory usage scales predictably with model size
  • Efficiency scores favor SmoothQuant for most model sizes
Figure 6: Spindle Plot Comparison: Multi-Metric Analysis Across Models and Methods
A.3.4 t-SNE Embedding Analysis
Figure 7 presents t-SNE embeddings of quantized weight distributions, revealing the structural similarities and differences between quantization methods. The analysis provides insights into:
  • Distribution Clustering: Methods with similar quantization strategies cluster together
  • Outlier Detection: Identification of quantization methods that produce unusual weight distributions
  • Structural Preservation: Analysis of how well each method preserves the original weight structure
The t-SNE visualization shows that:
  • SmoothQuant and SimQuant produce similar weight distributions (clustered together)
  • FP16 forms a distinct cluster representing the baseline distribution
  • AWQ and GPTQ show intermediate clustering patterns
  • ZeroQuant exhibits the most distinct distribution pattern
Figure 7: t-SNE Embedding: Quantized Weight Distribution Analysis
A.3.5 Scaling Curves Analysis
Figure 8 presents comprehensive scaling curves showing quantization effectiveness across different model sizes and context lengths. The four subplots analyze:
  1. 1.
    Throughput Scaling: Performance degradation with increasing model size
  2. 2.
    Memory Scaling: Memory usage growth with model size
  3. 3.
    Context Length Scaling: Perplexity changes with sequence length
  4. 4.
    Efficiency Scaling: Overall efficiency trends with model size
Key findings from the scaling analysis:
  • Linear Scaling: Quantization overhead remains constant relative to model size
  • Context Efficiency: SimQuant shows superior performance for long sequences (32K+ tokens)
  • Memory Scaling: Near-linear memory reduction with increasing model size
  • Efficiency Trends: SmoothQuant maintains the best efficiency across all scales
Figure 8: Scaling Curves: Comprehensive Analysis of Quantization Effectiveness Across Scales
A.4 Statistical Analysis and Significance Testing
A.4.1 Performance Significance Testing
We conduct statistical significance testing to validate the performance differences observed in our experiments. Using paired t-tests with Bonferroni correction for multiple comparisons, we analyze:
  • Perplexity Differences: Statistical significance of accuracy improvements
  • Throughput Differences: Significance of speed improvements
  • Memory Efficiency: Statistical validation of memory savings
The results show that LLMEasyQuant’s improvements are statistically significant (p ¡ 0.01) across all metrics and models, confirming the robustness of our findings.
A.4.2 Confidence Intervals and Error Analysis
We provide 95% confidence intervals for all performance metrics, ensuring the reliability of our experimental results. The error analysis reveals:
  • Low Variance: Consistent performance across multiple runs
  • Reliable Improvements: Confidence intervals confirm performance gains
  • Robust Results: Statistical validation across different experimental conditions
A.5 Computational Complexity Analysis
A.5.1 Algorithmic Complexity
We analyze the computational complexity of LLMEasyQuant’s quantization algorithms:
Theorem 5 (SmoothQuant Complexity).
Proof of Complexity.
The SmoothQuant algorithm performs the following operations:
  1. 1.
    Scale factor computation: $O(D)$ time, $O(D)$ space
  2. 2.
    Activation smoothing: $O(B\cdot D)$ time, $O(B\cdot D)$ space
  3. 3.
    Weight scaling: $O(D\cdot D^{\prime})$ time, $O(D\cdot D^{\prime})$ space
  4. 4.
    Quantization: $O(B\cdot D+D\cdot D^{\prime})$ time, $O(B\cdot D+D\cdot D^{\prime})$ space
The overall complexity is dominated by the matrix operations, resulting in $O(B\cdot D\cdot D^{\prime})$ time complexity and $O(B\cdot D+D\cdot D^{\prime})$ space complexity. ∎
A.5.2 Memory Hierarchy Optimization
We analyze the memory hierarchy optimization in LLMEasyQuant’s fused kernels:
Theorem 6 (Memory Bandwidth Optimization).
@@ A.5.2 Memory Hierarchy Optimization
Proof of Memory Optimization.
For separate operations:
  • Load FP16 weights: $2\times|W|$ bytes
  • Store quantized weights: $b/8\times|W|$ bytes
  • Load quantized weights for GEMM: $b/8\times|W|$ bytes
  • Total: $(2+2\times b/8)\times|W|$ bytes
For fused operation:
  • Load FP16 weights: $2\times|W|$ bytes
  • Store quantized weights: $b/8\times|W|$ bytes
  • Total: $(2+b/8)\times|W|$ bytes
@@
A.6 Error Propagation Analysis
A.6.1 Layer-wise Error Accumulation
We analyze how quantization errors propagate through transformer layers:
Theorem 7 (Error Accumulation Bound).
@@ A.6.1 Layer-wise Error Accumulation
Since $\epsilon_{\ell}\leq\epsilon$ for all layers, the accumulated error is $O(L\cdot\epsilon)$. ∎
A.7 Calibration Data Requirements Analysis
A.7.1 Minimum Calibration Set Size
We derive the minimum calibration data requirements for accurate quantization:
Theorem 8 (Calibration Data Requirements).
Proof of Calibration Requirements.
where $n$ is the sample size and $[a,b]$ is the range of the data.
Setting the right-hand side to $\delta$ and solving for $n$:
This comprehensive theoretical analysis provides the foundation for LLMEasyQuant’s practical implementation and validates its design choices across quantization methods, distributed execution, and hardware optimization. The detailed mathematical formulations ensure rigorous understanding of the quantization processes and their theoretical guarantees.
Lemma 3.
Proof.
Since AllGather is a deterministic collective operation, all devices will have identical $\delta_{\ell}^{\text{global}}$ and $z_{\ell}^{\text{global}}$ after synchronization. This ensures that quantized weights $\hat{W}_{\ell}$ are identical across all devices, maintaining model consistency. ∎
A.8 Memory Hierarchy Optimization Analysis
A.8.1 HBM-SRAM Transfer Optimization
Theorem 9.
Proof.
For separate operations:
  • Load FP16 weights: $2\times|W|$ bytes
  • Store quantized weights: $b/8\times|W|$ bytes
  • Load quantized weights for GEMM: $b/8\times|W|$ bytes
  • Total: $(2+2\times b/8)\times|W|$ bytes
For fused operation:
  • Load FP16 weights: $2\times|W|$ bytes
  • Store quantized weights: $b/8\times|W|$ bytes
  • Total: $(2+b/8)\times|W|$ bytes
A.9 Quantization Error Propagation Analysis
A.9.1 Layer-wise Error Accumulation
Theorem 10.
Proof.
Since $\epsilon_{\ell}\leq\epsilon$ for all layers, the accumulated error is $O(L\cdot\epsilon)$. ∎
A.10 Performance Analysis of Fused Kernels
A.10.1 CUDA Kernel Efficiency
@@
Proof.
The kernel design uses:
  • Persistent thread blocks for reduced kernel launch overhead
  • Cooperative warp-level reductions for scale computation
  • Shared memory tiling to minimize global memory access
  • Tensor Core utilization for INT8 GEMM operations
@@
However, due to register pressure and shared memory usage, practical occupancy is $\geq 75\%$, which is optimal for memory-bound operations. ∎
A.11 Calibration Data Requirements
A.11.1 Minimum Calibration Set Size

### v6

@@ B.6.1 Minimum Calibration Set Size
Since AllGather is a deterministic collective operation, all devices will have identical $\delta_{\ell}^{\text{global}}$ and $z_{\ell}^{\text{global}}$ after synchronization. This ensures that quantized weights $\hat{W}_{\ell}$ are identical across all devices, maintaining model consistency.
B.7 Memory Hierarchy Optimization Analysis
B.7.1 HBM-SRAM Transfer Optimization
Theorem B.35.
@@ B.7.1 HBM-SRAM Transfer Optimization
Proof B.36.
For separate operations, the memory bandwidth requirements include loading FP16 weights ($2\times|W|$ bytes), storing quantized weights ($b/8\times|W|$ bytes), and loading quantized weights for GEMM ($b/8\times|W|$ bytes), resulting in a total of $(2+2\times b/8)\times|W|$ bytes.
For fused operation, the memory bandwidth requirements include loading FP16 weights ($2\times|W|$ bytes) and storing quantized weights ($b/8\times|W|$ bytes), resulting in a total of $(2+b/8)\times|W|$ bytes.
@@
B.8 Quantization Error Propagation Analysis
B.8.1 Layer-wise Error Accumulation
Theorem B.37.
@@ B.8.1 Layer-wise Error Accumulation
Since $\epsilon_{\ell}\leq\epsilon$ for all layers, the accumulated error is $O(L\cdot\epsilon)$.
B.9 Performance Analysis of Fused Kernels
B.9.1 CUDA Kernel Efficiency
@@
Proof B.40.
The kernel design employs persistent thread blocks for reduced kernel launch overhead, cooperative warp-level reductions for scale computation, shared memory tiling to minimize global memory access, and Tensor Core utilization for INT8 GEMM operations.
@@
However, due to register pressure and shared memory usage, practical occupancy is $\geq 75\%$, which is optimal for memory-bound operations.
B.10 Calibration Data Requirements
B.10.1 Minimum Calibration Set Size